OpenStudy (anonymous):

50% of what number is 25 this would be 50 right? because 50% is half so 25+25 is 50...Is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

in general : 50% of x is meangs : 50x/100 in your specific case : 50x/100 = 25 x = 250/5 x = 50

OpenStudy (anonymous):

ahh I see thank you :D